# Towards Contextual Sensitive Data Detection

### v1

@@ 5 Experimental Setup
Here, we outline the implementation details, the LLMs used, baselines, datasets, and metrics to evaluate our proposed framework for contextually sensitive data detection. Code and datasets can be found at https://github.com/trl-lab/sensitive-data-detection.
5.1 Method implementation
5.1.1 Type-contextualization
@@
In the detection step, given the well-defined nature of PII types, we parse the problem as a multiclass classification task over a taxonomy of personal identifiable types (e.g., email_address, generic_identifier) derived from GDPR [gdpr] and UN OCHA guidelines [OCHA2021Guidelines] (see the code repository for a complete list). The LLM is prompted with a column name and five values, to return only the most appropriate type or None. We consider zero-shot and fine-tuned approaches. The zero-shot method leverages the model’s inherent ability to follow instructions, while the fine-tuned approach enhances performance by training the models on a synthetic dataset of 1,000 labeled columns. Fine-tuning is done using Low-Rank Adaptation [LORA21_HU] for parameter efficiency. The models are optimized with AdamW for a single epoch using an 80/20 train-validation split.
@@
In the reflection phase, the LLM re-evaluates each candidate column corresponding to types of interest within the context of the entire table. We provide the LLM with a markdown-formatted representation of the full table, including headers and five sample rows per column, along with the specific type detected for the target column and, if applicable, the entities of other columns which are detected as PII entities. The LLM is then instructed to assign one of three contextual sensitivity levels: non_sensitive, referring to information that cannot identify a person (e.g., aggregate data, organization names); moderate_sensitive, referring to data that could potentially identify a person when combined with other attributes (e.g., demographics, partial information); and high_sensitive, referring to data that definitively identifies a person (e.g., full name, email, national ID).
5.1.2 Domain-contextualization
@@
For the retrieval step, we aim to identify relevant heuristics indicating sensitivity rules from domain-specific data governance documents in order to integrate this external domain knowledge in the LLM’s context. In our experiments we focus on a case study on the Humanitarian Data Exchange (HDX) [hdx], hence on humanitarian datasets. The system retrieves relevant policy fragments from Information Sharing Protocols (ISPs) based on the dataset’s geographical context (e.g., country). We use GPT-4o-mini to extract the specific sensitivity rules from these documents (e.g., data types for low, moderate, high, severe sensitivity), which we store in a dictionary. If no country-specific ISP is identified, a default fallback template based on cross-country consensus is used.
@@ 5.2 Model Selection
We evaluate closed-source and open-source models, to understand trade-offs between these options. From closed-source options, we select the GPT-4o-mini [openai2024gpt4ocard] model for its good performance and cost trade-off. We focus on open-source LLMs that enable local deployment, for which we select the Gemma family (2 and 3) [gemma_2025, Gemma2Paper] and Qwen3 in 8B and 14B sizes [yang2025qwen3technicalreport]. Qwen3 14B was selected given the strong performance of its predecessor Qwen2.5’s exceptional performance on benchmarks such as MMLU Pro  [^1] https://huggingface.co/spaces/TIGER-Lab/MMLU-Pro, demonstrating strong reasoning capabilities and robust general understanding. Furthermore, Gemma2 9B and Gemma3 12B were included to strike a balance between computational efficiency and performance, and their support for quantized deployment. We also include versions of Qwen3 8B and Gemma2 9B that are fFgine-tuned on a synthetic dataset of 1K table columns hand-labeled with 27 PII types, indicated with “FT” suffix.
@@ 5.3 Baselines
We compare our methods with two common baselines for sensitive data detection. We include Google Cloud DLP [google_DLP], a rule-based system that combines built-in infoTypes, predefined detectors for common PII entities, which are associated custom-defined dictionaries and regular expressions. For our experiments, we configure the minimum likelihood threshold to “Likely”. The second baseline is Microsoft Presidio, an open-source tool that leverages regular expressions, pattern-based rules, context windows, and ML-based named entity recognition to detect PII types. To enable consistent evaluation, each Presidio type has been mapped to a corresponding category in our unified taxonomy (this mapping can be found in the code repository). Both systems output column-level predictions and are widely adopted in enterprise settings. However, neither tool supports contextual reflection or table-wide reasoning; any match to a known PII pattern is automatically treated as sensitive.
@@ 5.4 Datasets
  • Type contextualization with sensitive PII: We extracted 66 anonymized tables (2,061 columns) from the GitTables corpus [hulsebos2023gittables] representing a variety of real-world database schemas. Each column was manually annotated with (i) its PII entity type and (ii) its contextual sensitivity level (non-sensitive, moderate or highly sensitive).
  • Domain contextualization with humanitarian tables: To evaluate the retrieve-then-detect mechanism, we generated 9 synthetic tables with sensitive data closely resembling typical humanitarian data (e.g., for displacement tracking, public health monitoring) for different countries. Each table contains columns that could be potentially sensitive within country-specific Information Sharing Protocols (ISPs). Humanitarian data quality experts labeled each column with a sensitivity level (low, moderate, high, severe). We also add 14 real tables from the Humanitarian Data Exchange [^2] https://data.humdata.org/ that do not contain any sensitive data, to evaluate the false positive rate.
@@ 5.5 Metrics
We evaluate along three axes: PII detection, contextual reflection for PII, and non-PII sensitivity detection. To account for class imbalance, we report both weighted (support-weighted) and macro averages. Weighted averages reflect overall performance across examples, while macro averages treat each class equally and reduce dominance by frequent classes. Unless otherwise noted, we focus on weighted averages.
Performance is assessed at both the column and table level. At the column level (type contextualization), PII detection is treated as a multiclass task over 27 PII types (plus a “None” category), while sensitivity reflection reduces to a binary decision between sensitive (moderate/high) and non-sensitive (low). At the table level (domain contextualization), a table is classified as sensitive if it contains any moderate/high/severe sensitive information, resulting in a binary task (sensitive vs. not). Precision, recall, and F1 are used as evaluation metrics.
For labeling, sensitivity levels are binarized: low is treated as non-sensitive, while moderate, high, and severe are grouped as sensitive. This aligns with operational practice where any moderate or higher sensitivity triggers safeguards, consistent with UN OCHA guidance that such data warrants protective measures [unocha2025].
6 Results of contextual sensitive data detection
We present a comprehensive evaluation of our proposed framework for contextually sensitive data detection in tabular datasets, through type- and domain-contextualization. Type Contextualization experiments instrumented by the detect-then-reflect mechanism are focused on personal-identifiable information (PII) and Domain Contextualization instrumented by the retrieve-then-reflect mechanism which focus on less standard data semantics through a humanitarian case-study.
@@ 6 Results of contextual sensitive data detection
System / Model Without Reflection With Reflection
@@
Google DLP 0.531 0.628 0.576
Presidio 0.520 0.618 0.565
GPT-4o-mini 0.856 0.639 0.732 0.938 0.632 0.755
Gemma 2 9B 0.740 0.819 0.778 0.800 0.792 0.796
Gemma 3 12B 0.487 0.941 0.642 0.753 0.806 0.779
Qwen3 8B 0.742 0.868 0.800 0.749 0.868 0.804
Qwen3 14B 0.565 0.972 0.714 0.732 0.941 0.824
Qwen3 8B FT $\rightarrow$ GPT-4o-mini 0.902 0.861 0.881
Without reflection, LLMs mark any detected PII-type column as sensitive, leading to high recall but lower precision; Gemma 3 12B and Qwen3 14B reach a recall of 0.94 and 0.97 but precision of only 0.49 and 0.57, respectively. GPT-4o-mini achieves higher precision, 0.86, at lower recall of 0.64. With reflection using full-table context, substantially improves precision across models, e.g., Qwen3 14B increases to 0.73 precision with 0.94 recall, trading a small decrease in recall for a significant gain on precision. A modular pipeline combining Qwen3 8B for type detection with GPT-4o-mini for reflection achieves F1 0.88, but is closely matched by Qwen3 14B which yields higher recall (0.94 compared to 0.86). These results illustrate the stark difference compared to existing baselines reported in the upper left quadrant, which suffer from insufficient recall ánd precision. Comparing the detect-then-reflect mechanism powered by open-source LLMs with commonly used pattern-matching tools DLP and Presidio, we observe a significant improvement in both precision and recall, addressing the key limitations of existing tools for deployment in practice. Overall, the detect-then-reflect mechanism corrects the naive assumption that all columns with potentially sensitive types are sensitive, boosting precision while maintaining recall.
6.1.1 Ablation: reflection without detection.
@@
Table 2: Reflection-only performance without prior type detection step, for identifying contextually sensitive columns. Evaluated across models using precision, recall, and F1 scores.
@@
GPT-4o-mini 0.964 0.562 0.711
Gemma 2 9B 0.878 0.750 0.809
Qwen3 8B 0.735 0.781 0.758
Qwen3 14B 0.761 0.885 0.819
@@
Shifting towards contextual sensitive data which cannot be assessed through a fixed set of types, we evaluate the retrieve-then-detect mechanism on a combined set of synthetic and real humanitarian tables from the HDX platform. Without external domain knowledge, i.e. sensitivity heuristics obtained through context retrieval, all models exhibit a highly conservative bias, achieving perfect recall but all with precision lower than 0.56 (Table 3). Integrating retrieved domain-specific context, in this case we leverage information sharing protocols (ISP’s) for humanitarian data sharing, into the prompt consistently improves precision. GPT-4o-mini’s precision increases from 0.47 to 0.69 (F1=0.82), and Qwen3 14B’s from 0.56 to 0.64 (F1=0.78), both maintaining perfect recall (1).
@@
System / Model No Domain Knowledge With Domain Knowledge
@@
Qualitative analysis reveals that the retrieval mechanism enables models to produce interpretable justifications, thereby enhancing the transparency and auditability of the sensitivity assessments. The below excerpt shows a context-grounded LLM assessment on a table titled “Reported location on NFIs suppliers” with a description of the data at hand (1), along with the LLM’s sensitivity assessment (2), and explanation grounded in retrieved ISP rules (3).
1. The column ‘‘nfi_supplier_loc- value -from_Iran’’ indicates the origin of NFIs suppliers.
2. This information, even with country-level disaggregation, doesn’t reveal personal or sensitive data. It relates to the location of suppliers, which is generally not considered sensitive.
@@
3. It aligns with the ISP’s guideline on ‘‘Facility data (e.g., health, education, water points) at national or regional level, unless explicitly restricted by clusters.’’ Since this column is about locations of suppliers and not about individuals or specific operations, it is categorized as non-sensitive.
The model outputs containing sensitivity classifications and explanations were also reviewed by two humanitarian data quality experts, which considered context-grounded explanations highly informative for consistency and improving human-assessment. Table 3 summarizes results on 9 datasets synthesized based on humanitarian scope, and 14 real humanitarian datasets from HDX without sensitive data to evaluate false positives.
7 Contextual Sensitive Data: importance, practice, and consequences
With the emergence of public data sharing portals, and their adoption in critical areas such as the Humanitarian Data Exchange (HDX) for humanitarian aid coordination, it is key to protect sensitive datasets before they are automatically retrieved from the portal by malicious users or web crawlers for AI training datasets. Not doing so, induces significant risk to vulnerable populations. We reconsider and broaden the definition of what data is sensitive, and argue that sensitive data should be assessed contextually. This shift is critical as prior research for detecting and handling sensitive data typically focuses mainly on personal-level sensitive data. To broaden this practice, we present mechanisms for detecting sensitive data based on the principles of type contextualization and domain contextualization, which form a steppingstone towards this ambition.
On real-world humanitarian datasets, we show that more contextual methods lead to higher precision, hence less false positive alarms, while preserving coverage. These mechanisms can enable automation but also assist humans in data reviewing by providing context-grounded model explanations and concrete citations to relevant sensitivity rules in domain-specific documents. For successful adoption in practice, reasonable sensitivity instructions should be available in the form of 1) fine-grained semantic types of data that are deemed sensitive for type contextualization, and 2) reasonably specified assessment rules in domain-specific documents.
A potential negative consequence of automated data sensitivity detection is that critical datasets might be blocked that are time-sensitive and could be instrumental for solving societal problems, for example, in humanitarian domains. The mechanisms could be tailored around the desired precision/recall trade-off to reduce false alarms. At this moment, the methods are only developed for data sensitivity detection, and not for automated remedies. In practice, a report could be generated based on existing cases to instruct the user on how to remedy the sensitive data or potentially even automatically remedy this.

### v2

@@ 4 Experimental Setup
Here, we outline the method implementation details, baselines, datasets, and metrics to evaluate our proposed framework for contextually sensitive data detection. Code and datasets can be found at https://github.com/trl-lab/sensitive-data-detection.
4.1 Implementation of the Mechanisms
The core components of the mechanisms are type detection, reflection, and retrieval. Further details of the implementation of these components per mechanism are provided below.
Type-contextualization
@@ 4.1 Implementation of the Mechanisms
In the detection step, given the well-defined nature of PII types, we parse the problem as a multiclass classification task over a taxonomy of personal identifiable types (e.g., email_address, generic_identifier) derived from GDPR European Parliament and Council of the European Union (2016) and UN OCHA guidelines UN OCHA Centre for Humanitarian Data (2021) (see the code repository for a complete list). We devise a language model to classify the type of a table column (or return None) Kayali et al. (2024); Shen et al. (2025), by prompting it with the column name and five values. We consider zero-shot and fine-tuned approaches. The zero-shot method leverages the model’s inherent ability to follow instructions, while the fine-tuned approach enhances performance by training the models on a synthetic dataset of 1,000 labeled columns. Fine-tuning is done using Low-Rank Adaptation Hu et al. (2021) for parameter efficiency. The models are trained for a single epoch using an 80/20 train-validation split.
@@
In the reflection phase, the LLM re-evaluates each candidate column corresponding to types of interest within the context of the entire table. We provide the LLM with a markdown-formatted representation of the full table, including headers and five sample rows per column, along with the specific type detected for the target column and, if applicable, the types of other columns which are detected as PII types. The LLM is then instructed to assign one of three contextual sensitivity levels: non_sensitive, referring to information that cannot identify a person (e.g., aggregate data, organization names); moderate_sensitive, referring to data that could potentially identify a person when combined with other attributes (e.g., demographics, partial information); and high_sensitive, referring to data that definitively identifies a person (e.g., full name, email, national ID).
Domain-contextualization
@@
Given the domain-specific nature of this mechanism, we conduct a case study with humanitarian data experts from the UN OCHA Humanitarian Data Centre using humanitarian crisis datasets from the Humanitarian Data Exchange (HDX) United Nations (2014). For the retrieval step, we aim to identify relevant heuristics indicating sensitivity rules from domain-specific data governance documents. The system retrieves relevant policy fragments from Information Sharing Protocols (ISPs) based on the dataset’s geographical context (e.g., country). We use GPT-4o-mini to extract and index the specific sensitivity rules from these documents. If no region-specific ISP is identified, a default ISP template based on cross-country consensus is used.
@@ 4.2 Model Selection
We evaluate closed-source and open-source models to understand trade-offs between these options, if any. While one of the goals of this work is to prevent sensitive data from leaking into proprietary LLMs, these LLMs are typically available under privacy-preserving protocols at some cost which can then provide a viable low-maintenance solution. From closed-source options, we select the GPT-4o-mini (OpenAI et al., 2024) model for its good performance and cost trade-off. We focus on open-source LLMs that enable local deployment, for which we select the Gemma family (2 and 3) (Team, 2025; Team et al., 2024) and Qwen3 in 8B and 14B sizes (Yang et al., 2025). Qwen3 14B was selected given the strong performance of its predecessor Qwen2.5’s exceptional performance on benchmarks such as MMLU Pro  [^1] https://huggingface.co/spaces/TIGER-Lab/MMLU-Pro, demonstrating strong reasoning capabilities and robust general understanding. Furthermore, Gemma2 9B and Gemma3 12B were included to strike a balance between computational efficiency and performance, and their support for quantized deployment. We also include versions of Qwen3 8B and Gemma2 9B that are fine-tuned on a synthetic dataset of 1K table columns hand-labeled with 27 PII types, indicated with “FT” suffix.
@@ 4.3 Baselines
We compare our methods with two common baselines for sensitive data detection. We include Google Cloud DLP Google Cloud (2018), a rule-based system that combines built-in infoTypes, predefined detectors for common Personal Identifiable Information (PII) types, which are associated custom-defined dictionaries and regular expressions. For our experiments, we configure the minimum likelihood threshold to “Likely”. The second baseline is Microsoft Presidio, an open-source tool that leverages regular expressions, pattern-based rules, context windows, and ML-based named entity recognition to detect PII types. To enable consistent evaluation, each Presidio type has been mapped to a corresponding category in our unified taxonomy (this mapping can be found in the code repository). Both systems output column-level predictions and are widely adopted in enterprise settings. However, neither tool supports contextual reflection or table-wide reasoning; any match to a known PII pattern is automatically treated as sensitive.
@@ 4.4 Datasets
  • Type contextualization with sensitive PII: We extracted 66 anonymized tables, with in total 2,061 columns, from the GitTables corpus Hulsebos et al. (2023) representing a variety of real-world table schemas. Each column was manually annotated with (i) its PII type and (ii) its contextual sensitivity level (non-sensitive, moderate, or highly sensitive).
  • Domain contextualization with humanitarian data: To evaluate the retrieve-then-detect mechanism for domain-specific sensitive data, we conduct a case study with humanitarian experts from the UN OCHA Humanitarian Data Centre. We collect 23 tabular datasets, with in total 326 columns, resembling typical humanitarian datasets (e.g., for displacement tracking, public health monitoring) for different regions. With rich input from humanitarian experts to ensure realistic properties, we synthesize 9 tabular datasets with one or more actual contextually sensitive data in line with region-specific Information Sharing Protocols (ISPs). Two humanitarian data experts annotated columns (100 in total, of which 70 contain sensitive data) with a sensitivity level (low, moderate, high, severe). The other 14 tabular datasets are similar but real tabular datasets from the UN’s Humanitarian Data Exchange [^2] https://data.humdata.org/ and have been manually curated as challenging datasets without actual sensitive data, to evaluate the false positive rate.
Evaluation granularity
For evaluation, sensitivity levels are binarized: low is treated as non-sensitive, while moderate, high, and severe are grouped as sensitive. This aligns with operational practice in humanitarian contexts where any moderate or higher sensitivity triggers safeguards, consistent with UN OCHA guidance that such data warrants protective measures (OCHA, 2025). For domain contextualization, we assess performance on table-level to align with the ISP’s coarse-grained data policy instructions, despite having column-level annotations. Tables are labeled sensitive if one of its columns is annotated as sensitive (moderate, high or severe).
@@ 4.5 Metrics
We evaluate accuracy along three axes: PII detection, contextual reflection for PII, and non-PII sensitivity detection. To account for class imbalance, we report (support-)weighted averages. Performance is measured at column and table level. For type contextualization we evaluate column-level accuracy, PII detection is treated as a multiclass task over 27 PII types (plus a “None” category), while sensitivity reflection reduces to a binary decision between sensitive (moderate/high) and non-sensitive (low). For domain contextualization we evaluate table-level accuracy: a table is classified as sensitive if it contains any moderate/high/severe sensitive information, resulting in a binary task (sensitive vs. not). Precision, recall, and F1 are used as evaluation metrics.
Besides accuracy metrics, we report the average runtime per column to assess deployment considerations.
5 Results
We present a comprehensive evaluation of our proposed framework for contextually sensitive data detection in tabular datasets, through type- and domain-contextualization. Type Contextualization experiments instrumented by the detect-then-reflect mechanism are focused on personal-identifiable information (PII) and Domain Contextualization instrumented by the retrieve-then-detect mechanism which focuses on domain-specific sensitivity through a case-study with humanitarian data experts.
@@ 5 Results
system/model w/o reflection w/ reflection
@@
Google DLP 0.53 0.63 0.58
Presidio 0.52 0.62 0.57
GPT-4o-mini 0.86 0.64 0.73 0.94 0.63 0.76
Gemma 2 9B 0.74 0.82 0.78 0.80 0.79 0.80
Gemma 3 12B 0.49 0.94 0.64 0.75 0.81 0.78
Qwen3 8B 0.74 0.87 0.80 0.75 0.87 0.80
Qwen3 14B 0.57 0.97 0.71 0.73 0.94 0.82
Qwen3 8B FT -> GPT-4o-mini 0.90 0.86 0.88
Without reflection, LLMs mark any detected PII-type column as sensitive, leading to high recall but lower precision; Gemma 3 12B and Qwen3 14B reach a recall of 0.94 and 0.97 but precision of only 0.49 and 0.57 respectively (Table 1). GPT-4o-mini achieves higher precision, 0.86, at lower recall of 0.64. With reflection using full-table context, substantially improves precision across models, e.g., Qwen3 14B increases to 0.73 precision with 0.94 recall, trading a small decrease in recall for a significant gain on precision. A modular pipeline combining Qwen3 8B for type detection with GPT-4o-mini for reflection achieves F1 0.88, but is closely matched by Qwen3 14B which yields higher recall (0.94 compared to 0.86). These results illustrate the stark difference compared to existing baselines reported in the upper left quadrant, which suffer from insufficient recall ánd precision. Comparing the detect-then-reflect mechanism powered by open-source LLMs with commonly used pattern-matching tools DLP and Presidio, we observe a significant improvement in both precision and recall, addressing the key limitations of existing tools for deployment in practice. Overall, the detect-then-reflect mechanism corrects the naive assumption that all columns of sensitive types are indeed always sensitive, boosting precision while maintaining recall.
Ablation: reflection without detection.
@@
Table 2: Reflection-only performance without prior type detection, for identifying contextually sensitive columns. Evaluated across models using precision, recall, and F1 scores.
@@
GPT-4o-mini 0.96 0.56 0.71
Gemma 2 9B 0.88 0.75 0.81
Qwen3 8B 0.74 0.78 0.76
Qwen3 14B 0.76 0.89 0.82
@@
Shifting towards contextual sensitive data which cannot be assessed through a fixed set of types, we evaluate the retrieve-then-detect mechanism on a combined set of synthetic and real humanitarian tables from the HDX platform. Without external domain knowledge, all models exhibit a highly conservative bias, achieving perfect recall but all with precision lower than 0.56 (Table 3; baseline). Integrating retrieved domain-specific context, i.e., sensitivity guidelines obtained from information sharing protocols (ISP’s) for humanitarian data sharing, into the input for reflection by the language model, consistently improves precision. GPT-4o-mini’s precision increases from 0.47 to 0.69 (F1=0.82), and Qwen3 14B’s from 0.56 to 0.64 (F1=0.78), both maintaining perfect recall (1). Table 3 summarizes the results on the 23 datasets humanitarian datasets.
@@
system/model w/o domain knowledge w/ domain knowledge
@@
Qualitative analysis surfaces that the context retrieval mechanism enables models to produce interpretable and consistent justifications, thereby enhancing the transparency and auditability of the sensitivity assessments. The model outputs containing sensitivity classifications and explanations were reviewed by two humanitarian data quality experts, who considered the context-grounded explanations accurate, and highly informative for guiding and standardizing human-assessment. The below excerpt shows such context-grounded LLM assessment on a table titled “Reported location on NFIs suppliers” with (1) a description of the data at hand, along with (2) the LLM’s sensitivity assessment, and (3) explanation grounded in retrieved ISP rules.
@@
1. The column ‘‘nfi_supplier_loc-value-from_Iran’’ indicates the origin of NFIs suppliers. 2. This information, even with country-level disaggregation, doesn’t reveal personal or sensitive data. It relates to the location of suppliers, which is generally not considered sensitive. 3. It aligns with the ISP’s guideline on ‘‘Facility data (e.g., health, education, water points) at national or regional level, unless explicitly restricted by clusters.’’ Since this column is about locations of suppliers and not about individuals or specific operations, it is categorized as non-sensitive.
5.3 Error analysis
Type contextualization
On data with PII, we particularly observed misclassification of columns of generic types, as well as explicit personal types. For example, the “GENERIC_ID” type is well detected, but classified as non-sensitive by most LLMs causing very low recall. For types more explicitly relating to personal information (like names), columns are often classified as sensitive in reflection, while not always annotated as such. This highlights the potential for more specific context- and type-specific instructions during reflection.
Domain contextualization
In our experiments with humanitarian data, we observed mainly false positives for LLMs, which generally appear cautious with humanitarian data. On these false positives, analysis surfaces that the Qwen model has most confusion around borderline cases that are annotated as non-sensitive but classified as medium-sensitive, while gpt-4o-mini deems most of these tables highly sensitive (75%). This illustrates the need for more specific instructions for borderline cases.
5.4 Latency of contextual sensitive data detection.
Table 4 gives an overview of wall-clock latency measurements of the end-to-end pipeline averaged over all columns, as well as cost indications and memory measurements for local models. Despite the more elaborate mechanisms for contextualized sensitive data detection, the latency delta between the contextual sensitive LM-powered pipelines versus Google DLP (which only conducts type-based sensitivity detection) might be acceptable given the significant accuracy gain that the introduced mechanisms provide. If efficiency is key, smaller models like Qwen3 8B can yield an favorable trade-off, whereas larger or more costly models for more sensitive scenarios can be feasible for smaller-scale evaluations.
Table 4: Comparison of pipeline latency per column and resource cost across different methods and models.
tool/model total pipeline (s/col) cost or GPU memory
Google DLP [^3] For Google DLP, we only measure type-based sensitivity detection as it is not suitable for domain-specific sensitivity. 0.16 subscription
GPT-4o-mini 1.20 \$0.05 per table
Gemma 2 9B 2.89 7.8 GB
Gemma 3 12B 2.94 8.0 GB
Qwen3 8B 0.46 8.0 GB
Qwen3 14B 1.62 11.4 GB
6 Related Work
Existing methods for detecting sensitive data concentrate on detecting PII types, or otherwise static sensitive types. To this end, widely adopted industry tools like Microsoft Presidio (Microsoft, 2018) and Google Cloud DLP (Google Cloud, 2018) detect correspondences between values and structured types, such as email addresses and phone numbers, based on rule-based heuristics and syntactic pattern matching. Recent research has expanded the type detection scope by utilizing large neural models to learn mappings between high-dimensional semantic representations of values, e.g. table columns, and their semantic types Hulsebos et al. (2019); Kužina et al. (2023). Language models have also emerged as powerful tools for sensitive data detection, as particularly shown for PII detection, outperforming both rule-based and traditional machine learning methods in accuracy and adaptability (Shen et al., 2025; Yang et al., 2023). However, while language models excel at identifying known PII types, they struggle with assessing contextual sensitivity, such as domain-specific sensitive types, unless explicitly provided with the relevant operational context. In this work, we address this gap: our contextual sensitive data detection framework addresses both personal and non-personal sensitive data by considering internal and external context factors.